### v1

# Machine learning-based identification of Gaia astrometric exoplanet orbits

Abstract
The third Gaia data release (DR3) contains $\sim$170 000 astrometric orbit solutions of two-body systems located within $\sim$500 pc of the Sun. Determining component masses in these systems, in particular of stars hosting exoplanets, usually hinges on incorporating complementary observations in addition to the astrometry, e.g. spectroscopy and radial velocities. Several DR3 two-body systems with exoplanet, brown-dwarf, stellar, and black-hole components have been confirmed in this way. We developed an alternative machine learning approach that uses only the DR3 orbital solutions with the aim of identifying the best candidates for exoplanets and brown-dwarf companions. Based on confirmed substellar companions in the literature, we use semi-supervised anomaly detection methods in combination with extreme gradient boosting and random forest classifiers to determine likely low-mass outliers in the population of non-single sources. We employ and study feature importance to investigate the method’s plausibility and produced a list of 22 best candidates of which four are exoplanet candidates and another five are either very-massive brown dwarfs or very-low mass stars. Three candidates, including one initial exoplanet candidate, correspond to false-positive solutions where longer-period binary star motion was fitted with a biased shorter-period orbit. We highlight nine candidates with brown-dwarf companions for preferential follow-up. One candidate companion around the Sun-like star G 15-6 could be confirmed as a genuine brown dwarf using external radial-velocity data. This new approach is a powerful complement to the traditional identification methods for substellar companions among Gaia astrometric orbits. It is particularly relevant in the context of Gaia DR4 and its expected exoplanet discovery yield.
keywords:
planets and satellites: detection – brown dwarfs – binaries: general – astrometry – methods: data analysis – surveys
pubyear: 2024pagerange: Machine learning-based identification of Gaia astrometric exoplanet orbitsA
1 Introduction
The third Gaia data release (DR3, Gaia Collaboration et al., 2016, 2023a) delivered the first uniform and large-scale census of astrometric binaries and increased the number of available orbital solutions by two orders of magnitude from a few thousand (Mason et al., 2001) to $\sim$170 000. The revolutionary accuracy of Gaia made it possible to determine a large number of astrometric orbits with semi-major axes smaller than 1 milli-arcsecond (mas) for the first time (Halbwachs et al., 2023), finally opening the astrometric exoplanet-science window widely. However, small semi-major axes do not necessarily imply low companion masses: because the DR3 astrometric orbits usually refer to the unresolved photocentre of a binary, the principal astrophysical false-positive scenario for an exoplanet candidate is a nearly equal-mass and equal-brightness binary star (e.g. Gaia Collaboration et al., 2023b; Marcussen & Albrecht, 2023).
The astrometric orbits of several known giant exoplanets have been determined in DR3 (Holl et al., 2023; Gaia Collaboration et al., 2023b; Winn, 2022), leading to better constraints on their orbital configurations and masses. Gaia DR3 also contains a number of new giant exoplanet candidates from astrometry. [^1] The Gaia exoplanet candidate list is at https://www.cosmos.esa.int/web/gaia/exoplanets The default procedure to confirm those is to gather additional data, e.g. spectroscopy and precision radial-velocities (RV), that exclude any false-positive scenario.
Here we propose an intermediate step to identify the most promising exoplanet candidates that can be prioritised for follow-up. Our machine learning approach relies only on DR3 without the need for external datasets. Because of the very small number of confirmed exoplanets with DR3 orbits, which we need to ‘train‘ our models, we use as proof-of-concept the identification of substellar companions in general, i.e. both extrasolar planets and brown-dwarf companions. Since the mass distribution of substellar companions is structured but continuous (e.g. Grether & Lineweaver, 2006; Sahlmann et al., 2011), this does not imply a loss of applicability, but it conveniently doubles the size of the ‘training‘ sample.
Figure 1: Overview of the data and different partitionings used. Source preselection was performed as described in Section 2.2.
2 Gaia DR3 Data selection and augmentation
We queried the gaiadr3.nss_two_body_orbit table and selected all astrometric orbits, i.e. entries having a nss_solution_type of either ‘Orbital’, ‘AstroSpectroSB1’, ‘OrbitalAlternative*’, or ‘OrbitalTargetedSearch*’.
For 98 sources that have both ‘AstroSpectroSB1’ and ‘OrbitalTargetedSearch*’ solutions we only retained the ‘AstroSpectroSB1’ solutions. Two solutions were identified as false positives by the Gaia collaboration [^2] https://www.cosmos.esa.int/web/gaia/dr3-known-issues#FalsePositive and we discarded those. This left us with 169 127 orbit solutions for the same number of sources.
For these sources, we retrieved the data of the gaia_source table which contains magnitudes, colours, radial velocities, and other information obtained from the astrometric single-star solution (Lindegren et al., 2021a).
The DR3 table fields are described in the online documentation (van Leeuwen et al., 2022), but not all are relevant for the purpose of our study. The selection of relevant fields and the augmentation with additional fields is described below.
The gaia_source and nss_two_body_orbit fields that we used are listed in Tables 5 and 6, respectively. We augmented these tables with a few fields to make relevant source or orbit parameters directly accessible and interpretable. For the gaia_source table, we computed the absolute source $G$-band magnitude from the apparent magnitude phot_g_mean_mag and the parallax $\varpi_{\mathrm{ss}}$ from gaia_source without zero-point correction (Lindegren et al., 2021b).
For the nss_two_body_orbit table, we converted the Thiele-Innes coefficients ($A,B,F,G$) into the geometric orbital elements (the semimajor axis of the photocentre orbit $a_{0}$, the argument of periastron $\omega$, the longitude of the ascending node $\Omega$, and the orbital inclination $i$) using the linearised formulae described in Halbwachs et al. (2023); Gaia Collaboration et al. (2023b); Holl et al. (2023) and taking into account all parameter covariances using code implemented in pystrometry (Sahlmann, 2019). We also computed the astrometric mass function in solar masses as described in Halbwachs et al. (2023, Eq. 13):
$f_{M}=365.25^{2}\frac{a_{0}^{3}}{P^{2}\varpi_{\mathrm{nss}}^{3}},$ (1)
where $P$ is the orbital period in days and $\varpi_{\mathrm{nss}}$ is the parallax from nss_two_body_orbit in mas. Finally we joined both tables on the source identifier source_id.
We note that the nss_solution_type is disregarded in the machine learning analysis. This is intentional because we want to apply our approach to purely astrometric orbits with as little selection effects as possible. A nss_solution_type = ‘AstroSpectroSB1’, for instance, implies that a compatible orbit was detected in two Gaia instruments at the same time, and the necessary RV amplitude makes it unlikely that an exoplanet signature corresponds to a ‘AstroSpectroSB1’.
2.1 Data imputation
For some sources/solutions not all table fields have been filled, e.g. sources that are too faint to be observed with the Gaia Radial Velocity Spectrometer (RVS, Katz et al., 2023) do not have an entry in the radial_velocity field. We therefore filled these occurrences with appropriate numerical values. The affected fields and fill-values are listed in Table 1. The fill values were determined as the median of the remaining population, expect for radial_velocity and eccentricity_error where they were set to zero.
Table 1: Fill values used for data imputation. The table nss_tbo stands for nss_two_body_orbit, FV stands for fill value, and $N$ indicates the number of values filled.
Table Field FV Unit $N$
gaia_source radial_velocity 0 km/s 21460
gaia_source radial_velocity_error 2.41 km/s 21460
gaia_source phot_rp_mean_mag 12.67 mag 2
gaia_source phot_bp_mean_mag 13.70 mag 2
gaia_source bp_rp 0.99 mag 2
gaia_source phot_bp_rp_excess_factor 1.22 2
nss_tbo eccentricity_error 0 1979
2.2 Definition of labels
To assign labels to individual sources and orbital solutions, we used the binary_masses table (Gaia Collaboration et al., 2023b, Sect. 5) and the literature.
131 142 of our selected sources have entries in binary_masses. 105 of those have several mass estimates stemming with different binary_masses.combination_method. We removed those listed as ‘SB2’ and ‘Eclipsing’ which left us with 131 037 unique sources. The preselected sources label was assigned to 1838 sources with binary_masses.m2_lower $<0.08$ and the high-probability preselected sources label was assigned to sources with binary_masses.m2_upper $<0.08$ (cf. Stevenson et al., 2023). In the context, the preselected sources label refers to solutions that correspond to sources, which based on their entries in binary_masses could be substellar-companion candidates.
Finally, we retrieved from the literature all the solutions that correspond to confirmed exoplanets (companion mass $M_{2}\lesssim 20M_{\mathrm{Jup}}$), brown-dwarf companions ($20M_{\mathrm{Jup}}\lesssim M_{2}\lesssim 80M_{\mathrm{Jup}}$), and binary stars ($0.12M_{\sun}\lesssim M_{2}$), where within that last category we also marked very-low-mass stellar companions with $80M_{\mathrm{Jup}}\lesssim M_{2}\lesssim 0.12M_{\sun}$. Table 2 shows the label counts and Table 7 lists the individual solutions, their labels, and the respective references.
Table 2: Label counts for our dataset.
Label Count
preselected sources 1787
high-probability preselected sources 26
brown_dwarf_companion 14
exoplanet 10
very_low_mass_stellar_companion 3
binary_star 3
false_positive_orbit 2
3 Data analysis
Figure 2: Overview of the algorithm used for identifying candidates.
3.1 Substellar Companions as Anomalies
Fundamentally, we expect the number of substellar companions in the non-single sources to be quite small, in the range of less than one percent and in the single digit percentile range amongst the preselected sources. As such, we can consider them to be outliers in our data, as can be appreciated in Figure 3, and we expect them to be identifiable based on their DR3 parameters.
Even though some relations between the astrometric data, such as the value of the mass function $f_{M}$ (see Equation 1), can be strongly indicative of the potential presence of a substellar companion, it is conceivable that there are various correlations and interactions between different parameters and the likelihood of a specific source having a substellar companion can likely not easily be reduced to a single parameter. Conversely, there are likely also other forms of outliers in the data, such as anomalies in the data processing pipeline or very low-mass stellar companions. Thus, we formulate the problem of identifying likely exoplanets and brown-dwarf companions as two-fold: first, in identifying outliers in the dataset and, secondly, identifying those outliers that are particularly likely to be substellar companions.
3.2 Semi-supervised Outlier Detection
Given the availability of some but few labelled examples (24 confirmed substellar companions), we choose a semi-supervised machine learning approach, i.e. one that relies on a small number of labelled samples and a larger number of unlabelled samples during training. Further, given the difficulty of solving the task at all due to the limited amount of confirmed samples we consider different dataset partitionings in the training and ensemble across configurations as later explained in this section.
In particular, as shown in Figure 1, we build four dataset versions (D1, D2, D3, D4), which contain subsets of the data (e.g. only preselected sources in D3 and D4) and add 26 high-probability preselected sources as positively labelled outliers (D2 and D4).
In most established machine learning approaches (Zhao & Hryniewicki, 2018; Han et al., 2022), the task of semi-supervised outlier detection is approached by training on the limited amount of data containing both known outliers and inliers and then predicting on a separate dataset. Since we only have a few confirmed outliers and virtually no confirmed inliers (i.e. binaries and very low mass stellar companions) we developed a different approach and instead consider our entire dataset of inliers (i.e. the vast majority of the data) to be weakly labelled, i.e. potentially mislabelled. Thus, we transform the task from classifying newly unseen samples into inliers or outliers into identifying the likeliest mislabelled samples in our training data. To this end, we perform a k-fold cross-validation and use the predictions on the presumed inliers as the indication of whether we believe that sample to be a mislabelled one. This procedure is shown in detail in Figure 2. We choose the amount of folds to be identical to the number of outliers to maximise the amount of confirmed outliers in the training data. The left out outlier is used to compute the validation metrics in the results section.
To further improve the robustness of the approach and since precise performance evaluation is challenging with the limited amount of data , we use two different approaches in our ensembles. In particular, we utilise XGBoost (Chen & Guestrin, 2016) as implemented in the Python package of the same name [^3] https://xgboost.readthedocs.io/en/stable/index.html Accessed: 2024-03-08 and a balanced random forest classifier (Chen et al., 2004) as implemented in the imblearn [^4] https://imbalanced-learn.org/ Accessed: 2024-03-08 Python module. Given the invariance of both approaches to monotonous transforms, no data transforms were applied beyond the elimination of samples (i.e. sources/orbital solutions) missing any of the studied features.
3.3 Feature Selection and Importance
The original data contain a total of 47 unique data points per source (an overview of all parameters is given in Tables 5 and 6). During the development of the approach we used different feature-importance measures such as SHapley Additive exPlanations (SHAP) values (Lundberg & Lee, 2017) and the feature importance implemented in scikit-learn (Pedregosa et al., 2011). Using this we identified and discarded several confounders, i.e. features that are marked as important for identifying outliers but that are not expected as such based on physical arguments.
As examples, we discarded both the parallax and the apparent magnitude of the source. These correlate positively with having confirmed substellar companions because those are the best-surveyed sources with ground-based, high-precision radial-velocity observations and generally are easier to follow-up. From an statistical astrophysical standpoint, there is little reason for a bright source to be more likely to host a substellar companion than a faint one, at least within the parameter space surveyed here. The other confounders that we identified are indicated in Tables 5 and 6.
Additionally, we use SHAP values as provided by the Python module [^5] https://shap.readthedocs.io/en/latest/ Accessed: 2024-03-08 of the same name to study the feature importance amongst our proposed candidates as well as to identify mislabelled training samples as shown in Section 4.
3.4 Ensembling and Candidate Selection
As outlined in the previous paragraphs, we use four different datasets (see Figure 1) and two different models (XGBoost and BalancedRandomForest). From these, we build our predictions using max-voting on the 50 samples with the highest outlier prediction as computed using the algorithm illustrated in Figure 2. We weight the voting relative to the chance to occur. Individual samples may, respectively, occur as inliers zero (confirmed substellar companions), two (high-probability preselected sources, orange in Figure 1), four (non-single sources) or eight (preselected sources) times. We measure the actual occurrence with three values: $\rho$ which describes the rate of occurrence in the eight runs and $\rho_{nss}$ and $\rho_{ssc}$, which are the rates of occurrence in the large datasets (D1 & D2) and smaller ones (D3 & D4), respectively.
4 Results: substellar-companion candidates
We applied a cutoff in relative occurrence at $\rho>0.5$ which resulted in the identification of our 14 best candidates for exoplanets and brown-dwarf companions. In addition we identified four candidates with $\rho_{nss}>0.5$ and four candidates with $\rho_{ssc}>0.5$. These three categories are presented in their respective sections in Table 4. All 22 candidates are shown in Figure 3 and discussed below. We also identified lower-confidence candidates using a lower cutoff value of $\rho>0.125$, which we list in Table 8.
Refer to captionRefer to caption
Figure 3: Density histograms of absolute magnitude as a function of colour (top panel, the inset show a zoom into the region of interest) and mass function (bottom) for all DR3 astrometric orbits. The blue contours indicate the concentration of the solutions labelled as preselected sources. Circles indicate the 10 confirmed exoplanets, squares indicate the 14 confirmed BD-companions, and crosses indicate our 22 best candidates for substellar companions discussed in the text. For reference, the Sun has $M_{G}=4.67$ and $G_{\mathrm{BP}}-G_{\mathrm{RP}}=0.82$ (Casagrande & VandenBerg, 2018), and the mass function of a $5\,M_{J}$ planet in a Jupiter-like orbit around the Sun is $1.1\cdot 10^{-7}$. These reference locations are marked with a yellow diamond. The exoplanet with the largest mass function is HD 39392 b (Wilson et al., 2016; Stevenson et al., 2023).
4.1 Validation
Ensuring the performance of our approach is a challenging task given the small amount of labelled samples. In the prediction method described in Figure 2 we obtain the predictions on the left-out fold as in a stratified K-fold validation. We use these predictions to compute performance metrics. As the inlier labels in these folds are weak labels (we expect some substellar companions in them) the metrics are probably underestimating the true performance since the models correctly "mislabel" some false inliers. Similarly, for the high probability (but not confirmed) outliers some may be mislabelled outliers, further impacting the performance measurements. Detailed results for different configurations are given in Table 3. Overall, we achieve very good results on the area under the receiver operating characteristic (AUROC) and accuracy of predictions. The area under the precision recall curve (AUPRC) results are worse but this is expected given the weak labels and significant class imbalance (less than 0.03% labelled outliers amongst all sources).
Table 3: Validation results for different configurations showing accuracy of prediction as well as area under the receiver operating characteristic (AUROC) and area under the precision recall curve (AUPRC).
Metric Only Confirmed Outliers High Probability Outliers All Sources Candidate Sources XGBoost Random Forest All
Accuracy 0.97 0.94 0.98 0.96 0.99 0.92 0.96
AUROC 0.98 0.97 0.99 0.97 0.98 0.97 0.97
AUPRC 0.65 0.49 0.5 0.57 0.66 0.47 0.56
A secondary measure of the performance of our approach is the accuracy of the applied majority voting scheme to identify the best candidates from the eight configurations. A total of 170 unique sources show up in the top 50 outliers of the eight approaches (i.e. amongst all samples including presumed inliers and outliers). Among those 170, we find 22 of our 24 confirmed substellar companions among the 61 highest ranked samples across all configurations, each being identified in at least two separate runs. The only missing ones are the brown dwarf companions in LHS 1610 (Gaia DR3 43574131143039104, Fitzmaurice et al., 2023) and the brown dwarf in HD 89707 (Gaia DR3 3751763647996317056, Gaia Collaboration et al., 2023b). The one around LHS 1610 was likely missed due to the very-red colour of its M5-dwarf host star, which is uncommon in the confirmed substellar companions. The one around HD 89707 was barely below the detection threshold scoring highly on most models. We investigate this further in Section 4.4 on feature importance.
Table 4: Best candidates for exoplanets and brown-dwarf companions. The SpT column indicates the source’s spectral type as listed in Simbad, the solution type column indicates a short-hand for the nss_solution_type (similar to Marcussen & Albrecht 2023: OTSV is ‘OrbitalTargetedSearchValidated’, OTS is ‘OrbitalTargetedSearch’, and ASB1 is ‘AstroSpectroSB1’), $M_{2}$ is a companion-mass estimate under the dark-companion assumption and using the primary-mass estimate $M_{1}$ from binary_masses, $\rho$ is the relative occurrence introduced in Section 3.4, $\rho_{\mathrm{nss}}$ is the relative occurrence for dataset variants D1&D2 only, $\rho_{\mathrm{ssc}}$ is the relative occurrence for dataset variants D3&D4 only, the $ssc$ flag is 1 when the solution was part of ‘preselected sources ‘ or ‘high-probability preselected sources ‘ and 0 otherwise, and $M_{2,\mathrm{alt}}$ is the companion-mass range as estimated in binary_masses.
Gaia DR3 source_id Name SpT sol. $P$ $f_{M}$ $M_{2}$ $\rho$ $\rho_{\mathrm{ssc}}$ $\rho_{\mathrm{nss}}$ $ssc$ $M_{1}$ $M_{2,\mathrm{alt}}$
type (day) ($M_{\sun}$) ($M_{\mathrm{Jup}}$) ($M_{\sun}$) ($M_{\sun}$)
2884087104955208064 HD 40503 K2V OTSV 826.5 1.9e-07 5.2 0.8 0.8 0.8 1 ${0.79}^{+0.05}_{-0.23}$ $0.004$-$0.681$
522135261462534528 * 54 Cas F8 OTS 401.1 2.5e-07 7.1 0.9 1.0 0.8 1 ${1.09}^{+0.06}_{-0.31}$ $0.005$-$0.905$
1712614124767394816 BD+75 510 M0V OTSV 297.6 6.4e-07 7.3 0.8 0.8 0.8 1 ${0.71}^{+0.05}_{-0.20}$ $0.006$-$0.632$
1610837178107032192 HD 128717 F8 OTS 1089.2 1.4e-06 12.8 0.9 1.0 0.8 1 ${1.13}^{+0.06}_{-0.16}$ $0.010$-$1.109$
1878822452815621120 BD+24 4592 K2 OTS 1009.3 5.4e-06 16.3 0.9 1.0 0.8 1 ${0.82}^{+0.05}_{-0.21}$ $0.014$-$0.766$
1897143408911208832 HD 207740 G5V OTS 712.2 2.8e-05 32.3 0.9 1.0 0.8 1 ${0.99}^{+0.05}_{-0.12}$ $0.029$-$1.089$
6330529666839726592 TYC 4998-437-1 Orbital 526.1 7.1e-05 39.0 0.6 0.5 0.8 1 ${0.82}^{+0.05}_{-0.20}$ $0.032$-$0.831$
364792020789523584 BD+32 92 F8 OTS 291.9 1.6e-04 67.2 1.0 1.0 1.0 1 ${1.21}^{+0.06}_{-0.14}$ $0.057$-$1.382$
5773484949857279104 GSC 09436-01089 ASB1 409.9 2.3e-04 73.8 0.6 0.8 0.5 1 ${1.16}^{+0.07}_{-0.12}$ $0.063$-$1.391$
3913728032959687424 HD 99251 F5 Orbital 328.9 2.3e-04 78.6 0.9 1.0 0.8 1 ${1.28}^{+0.06}_{-0.20}$ $0.069$-$1.411$
2540855308890440064 N/A Orbital 239.8 2.8e-04 65.7 0.9 1.0 0.8 1 ${0.88}^{+0.06}_{-0.17}$ $0.056$-$1.006$
1156378820136922880 G 15-6 G4 Orbital 268.3 3.6e-04 72.2 0.8 1.0 0.5 1 ${0.88}^{+0.05}_{-0.30}$ $0.065$-$0.872$
2171489736355655680 HD 206484 F8 ASB1 479.5 4.0e-04 90.4 0.8 1.0 0.5 1 ${1.18}^{+0.06}_{-0.06}$ ${0.078}^{+0.006}_{-0.006}$
3909531609393458688 TYC 277-599-1 Orbital 520.6 4.4e-04 79.5 0.8 1.0 0.5 1 ${0.92}^{+0.06}_{-0.13}$ $0.067$-$1.141$
4545802186476906880 HD 153376 F8V OTS 910.8 1.2e-05 24.2 0.4 0.0 0.8 0 NaN NaN
5148853253106611200 LP 769-9 Orbital 339.6 1.5e-04 44.8 0.4 0.0 0.8 1 ${0.69}^{+0.05}_{-0.19}$ $0.039$-$0.723$
3921176983720146560 HD 106888 F8 OTSV 366.3 5.8e-04 98.9 0.5 0.0 1.0 0 ${1.11}^{+0.06}_{-0.14}$ $0.083$-$1.388$
3067074530201582336 BD-06 2423A OTS 43.6 2.5e-03 155.8 0.4 0.0 0.8 0 NaN NaN
5484481960625470336 HD 49264 G3V Orbital 428.0 1.6e-04 59.7 0.5 0.8 0.2 1 ${1.02}^{+0.06}_{-0.08}$ $0.052$-$1.261$
2280560705703031552 HD 212620 F8 Orbital 429.0 2.0e-04 76.1 0.4 0.8 0.0 1 ${1.32}^{+0.06}_{-0.19}$ $0.065$-$1.475$
5323844651848467968 HD 78631 F8V ASB1 666.5 1.1e-03 123.6 0.5 1.0 0.0 1 ${1.11}^{+0.06}_{-0.06}$ ${0.076}^{+0.014}_{-0.012}$
1576108450508750208 HD 104289 F8 OTS 1233.3 1.2e-03 132.7 0.5 1.0 0.0 1 ${1.15}^{+0.06}_{-0.05}$ ${0.047}^{+0.006}_{-0.006}$
4.2 Exoplanet candidates
Table 4 is ordered by increasing mass function and we will discuss the individual systems here. We identified five high-probability exoplanet candidates. All of them are previously-known exoplanet candidates and have Gaia astrometric object of interest identifiers (Gaia-ASOI-# [^6] https://www.cosmos.esa.int/web/gaia/exoplanets):
The companion of HD 40503 (Gaia-ASOI-002, Gaia Collaboration et al., 2023b; Holl et al., 2023; Marcussen & Albrecht, 2023; Winn, 2022) continues to await full RV confirmation. The companion of * 54 Cas (Gaia-ASOI-003) is an exoplanet candidate discussed in Gaia Collaboration et al. (2023b). The star BD$+$75 510 (HIP 66071, Gaia-ASOI-001, Holl et al., 2023; Gaia Collaboration et al., 2023b) was followed up with high-precision RVs (Sozzetti et al., 2023) and a sub-Jupiter mass exoplanet was discovered. We also identify HD 128717 (Gaia-ASOI-009) as a strong candidate for hosting a super-Jupiter, motivating prioritised follow-up of this star.
The companion of BD$+$24 4592 (Gaia-ASOI-005) appears to be a strong exoplanet candidate in our analysis on the basis of the Gaia DR3 orbit, but this star was previously identified as spectroscopic binary (G 127--33, Latham et al., 2002) with a period of 3977 days and it is listed in the SB9 catalog (Pourbaix et al., 2004) with a large RV amplitude of 2.7 km/s. The Gaia astrometric period of 1009 days is very close to the DR3 time range. It is therefore plausible that the Gaia orbit is spurious and the partially-covered binary-star orbit was interpreted as a low-amplitude orbit at $\sim$1/4 of the true period. We therefore argue that Gaia-ASOI-005 should be re-classified as questionable.
4.3 Brown-dwarf companion candidates
We identified 17 brown-dwarf companion candidates with mass estimates spanning the brown-dwarf desert but predominantly located on the heavy side ($\gtrsim 60\,M_{J}$).
We identified the stars TYC 4998-437-1, BD+32 92, TYC 277-599-1, LP 769-9, HD 49264, HD 212620 as high-probability candidates for hosting brown-dwarf companions. For these six sources we did not find auxiliary information in the literature.
HD 207740 hosts a brown-dwarf companion candidate. Three Keck HIRES spectra of this source are discussed in Rice & Brewer (2020) but no RVs appear to have been published.
The stars HD 206484 and HD 78631 have DR3 solution types of ‘AstroSpectroSB1’ which means that the Gaia RVS also detected a compatible RV variation. The companion mass estimates from binary_masses in these cases are therefore more reliable and straddle the substellar boundary, i.e. these companions are either high-mass brown dwarfs or very-low-mass stars. For HD 78631 we noted a strong discrepancy between our companion-mass estimate ($\sim$123 $M_{\mathrm{Jup}}$, assuming a dark companion) and the one from binary_masses of $\sim$80 $M_{\mathrm{Jup}}$, which was determined using a non-zero flux ratio (Gaia Collaboration et al., 2023b).
The star HD 99251 hosts a brown-dwarf companion candidate. This star is also EPIC 201928968, a Kepler eclipsing binary candidate with orbital period $<$1 day. No close companion was found with Speckle observations (Schmitt et al., 2016).
The source Gaia DR3 2540855308890440064 hosts a brown-dwarf companion candidate. This source_id is not resolved by Simbad and the source is a 10″ visual binary with TYC 4672-237-1.
[…]
GSC 09436-01089 has a solution type of ‘AstroSpectroSB1’ which means that the Gaia RVS also detected a compatible RV variation. Using the primary-mass estimate of ${M_{1}}=1.16\pm 0.12\,M_{\sun}$ from binary_masses and resampling the DR3 orbit parameters while accounting for all covariances we obtain a companion-mass estimate of ${M_{2}}={74.88}^{+9.20}_{-8.07}\,{M_{\mathrm{Jup}}}$. GSC 09436-01089 B is therefore likely a high-mass brown dwarf but it may also be a very-low-mass star.
[…]
Finally, we present G 15-6 as the first independently-validated substellar companion identified by our work. The DR3 orbit implies a massive BD companion in a 268 day orbit. This source had already been identified as a possible brown-dwarf host from RV monitoring by Latham et al. (2002) and it is listed in the SB9 catalogue with a period of 266 days. The matching periods and companion-mass estimates leave little doubt that the same orbital motion is detected both by Gaia and the RVs from Latham et al. (2002). Using the primary mass estimate of ${M_{1}}=0.67\,M_{\sun}$ from Latham et al. (2002) with a 10 % uncertainty, and resampling the DR3 orbit parameters while accounting for all covariances we obtain a companion-mass estimate of ${M_{2}}={62.43}^{+6.16}_{-5.44}\,{M_{\mathrm{Jup}}}$. This companion therefore lies firmly in the substellar range and is the first confirmed brown-dwarf companion that was identified with our machine learning approach.
4.4 Feature Importance Results
In these section, we study the model predictions using SHAP values (Lundberg & Lee, 2017) to identify both the reasons for the models’ predictions on the entire sample population and our candidate pool, but also for individual samples such as the confirmed brown dwarf the models missed.
4.4.1 Discussion of selected features
It is helpful to inspect the information content of a few features in more detail:
Mass function ($f_{M}$): A small mass function is a necessary condition for a system with a substellar companion [^7] We do not discuss brown-dwarf binaries here., but it is not a sufficient condition because twin-binary stars also have very small astrometric mass functions.
Radial-velocity error: As explained in the online documentation and Katz et al. (2023); Babusiaux et al. (2023), the RV error given in gaia_source has a stellar magnitude-dependent meaning. For bright sources (phot_g_mean_mag $\lesssim$13) it is derived from the uncertainty of the median of the RV timeseries. In that case, a large RV error can therefore stem from large individual RV uncertainties or from unmodelled RV variations, e.g. as present for a binary star. For faint sources (phot_g_mean_mag $\gtrsim$13), the RV error is determined by stacking the individual cross-correlation functions and therefore the sensitivity to binary-induced motion is different (cf. Andrew et al., 2022). In the sample of $\sim$ 170 000 astrometric orbits with available RV error, about 43 % fall in the ‘bright‘ category, and this percentage decreases to 36 % for the preselected sources. In addition to the bi-modal determination of the RV error, there is the general dependency on apparent stellar magnitude: the brighter the star, the higher the signal-to-noise in the spectra, and the smaller the RV error. As a result of these multiple dependencies it is often not directly interpretable how the RV error has influenced the models’ predictions.
Goodness-of-fit: The goodness_of_fit from nss_two_body_orbit corresponds to the ‘gaussianized chi-square‘ statistic. For good orbital fits it should approximately follow a normal distribution with zero mean value and unit standard deviation. Values larger than roughly +3 indicate a bad fit. Our selection on astrometric orbits contains a mix of purely-astrometric orbits and combined RV+astrometric orbits, i.e. of type ‘AstroSpectroSB1’. For the latter, the goodness_of_fit refers to the combined fit, hence we can expect different feature characteristics between both groups.
Flux excess: the phot_bp_rp_excess_factor (Riello et al., 2021) contains a measure of the consistency between the red and blue source fluxes measured by Gaia. It has been shown that this metric can be sensitive to near-equal-brightness ($\Delta$mag $<0.5$) binaries (Belokurov et al., 2020).
Colour and absolute magnitude: Finally, the Gaia colour $G_{\mathrm{BP}}-G_{\mathrm{RP}}$ ($\equiv$ bp_rp) and absolute magnitude $M_{G}$ ($\equiv$ absolute_phot_g_mean_mag) indicate the location of the source in the colour-magnitude diagram (Fig. 3 top).
4.4.2 Predictions for populations
[…]
A detailed look at individual SHAP values in association with mass_function_msun values is available in Figure 6 which highlights their complex interaction. Conversely, the random sample in Fig. 5 displays that most samples are ruled out as candidates by large mass_function_msun values and/or large radial velocity errors. The Gaia colour bp_rp of the source also seems to play a more important role here. Note, however the difference in magnitude on the x-axes of Figures 4 and 5. Additionally, Figure 6 also displays the distributions of SHAP values and mass_function_msun values in random samples of non-single source population and the preselected sources. Note that even though they feature small mass_function_msun values, these are not associated with large SHAP values in most preselected sources.
Figure 4: SHAP value distribution for the top five features of the top 50 sources identified in each of the 8 configurations (total number of 8$\times$50 sources); colour indicates distribution of the feature; especially smaller values of the variables are associated with higher SHAP values.
Figure 5: SHAP value distribution for the top five features of in a random sample of 50 sources per configuration; colour indicates distribution of the feature; especially larger values of the variables are associated with smaller SHAP values.
Figure 6: SHAP values for mass_function_msun of candidates and random samples (same samples as in Fig. 5) in relation to their mass_function_msun values and in the background a histogram of the value distribution in the non-single source population
4.4.3 Predictions for individual systems
Investigating individual sources/solutions, we want to highlight three specific samples: The missed confirmed brown dwarf in LHS 1610, the top candidate around HD 40503, and Gaia DR3 246890014559489792 (Gaia object of interest identifier Gaia-ASOI-016), a small mass_function_msun outlier that was not identified as a candidate by our analysis. This can help us understand the decision process of the models, find potential confounders and help interpret the data of individual samples. In addition, decision plots for the missed brown dwarf around HD 89707 are in the Appendix in Figures 14 and 15. It was likely just barely below the detection threshold scoring highly on multiple models.
Figures 7 and 8 show decision plots for the missed brown dwarf around LHS 1610 for the random forest and XGBoost models, respectively. Here, the high radial-velocity error and mass_function_msun compared to other candidates such as HD 40503 are the main contributors to the decision. However, we can also see that the colour (bp_rp = 3.2) is a strong contributor, which makes sense as this is one of only two confirmed outliers with such a red colour - all others have bp_rp $<$ 1.5 as seen in Figure 3. Overall, we can indeed see that with those characteristics the brown dwarf around LHS 1610 is atypical compared to the other confirmed brown dwarfs and exoplanets and the model’s decision process is stringent. A remedy to this mis-identification would be the inclusion of more labelled training samples like the confirmed brown dwarf around LHS 1610.
Figures 9 and 10 display the decision plots for the top exoplanet candidate around HD 40503 for the random forest and XGBoost models, respectively. The classification as outlier is driven by small values of the radial-velocity error and mass_function_msun. Notably, the bp_rp colour was not an important feature.
Finally, Figures 11 and 12 show the decision plots for Gaia DR3 246890014559489792. Looking at the small mass_function_msun value of 1.367e-6 (cf. Figure 3) it would be conceivable that this is a strong candidate. However, as seen in the decision plots this is not the case because the other features, in particular the radial-velocity error, bp_rp, and goodness-of-fit contribute to an inlier classification.
Figure 7: Decision plot showing feature contribution to model output for the random forest models for the brown dwarf in LHS 1610. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 8: Decision plot showing feature contribution to model output for the XGBoost models for the brown dwarf in LHS 1610. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 9: Decision plot showing feature contribution to model output for the random forest models for HD 40503. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 10: Decision plot showing feature contribution to model output for the XGBoost models for HD 40503. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 11: Decision plot showing feature contribution to model output for the random forest models for GaiaDR3 246890014559489792. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 12: Decision plot showing feature contribution to model output for the XGBoost models for GaiaDR3 246890014559489792. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Additional data and a full list of candidates are given in the Appendix.
5 Discussion
The results shown in this work clearly indicate the viability of using a machine learning-based approach to identifying likely candidates for exoplanets and brown-dwarf companions. The majority of known exoplanets and brown-dwarf companions in the dataset were picked up successfully by the models and the identified new candidates exhibit physical properties, such as expected companion mass, that indicate that they indeed are worthy of follow-up studies.
One predictable drawback of the approach in this regard is the tendency of identifying new candidates that mostly lie within the distribution of already-known candidates in terms of their features. For example, the location of our best candidates in the colour-magnitude diagram (Fig. 3 top) largely coincides with the region where the confirmed substellar companions, i.e. the training sample, resides. This is a common drawback of machine learning approaches and could be remedied by providing a richer ‘training‘ dataset that, e.g., includes more confirmed substellar companions around redder sources (bp_rp > 1.5). In the same vein it has to be mentioned that the presented approach is of course greatly impacted by the very limited number of confirmed samples. These aspects will improve with time as the results of various DR3-orbit follow-up studies are being published and certainly with the release of Gaia DR4, when many more orbits of stars with substellar companions will be released. We anticipate that the iterative identification and confirmation of an increasing number of exoplanets with Gaia orbits, covering growing areas in parameter space, will lead to a positive feedback loop with machine learning approaches, which will help to accelerate the pace of discoveries.
A particularly precious dataset for approaches like ours are small mass-function systems that actually are confirmed to be near-twin binaries (Marcussen & Albrecht, 2023). These systems are crucial for teaching the models how to identify these astrophysical false positives. We therefore urge the teams involved in the observational follow-up of Gaia astrometric exoplanet candidates to publish their results, regardless of the outcome.
In terms of the performance and power of the presented approach several observations can be made. First, as displayed in Figures 4 and 5, the decision making of the models is truly multivariate as different features can lead to the identification of likely outliers and different features can support either a classification as outlier or not. For example, the goodness-of-fit feature is the third strongest in identification of our candidates but does not appear among the random sample were instead bp_rp is third strongest.
Another indicator of such behavior can be seen in Figure 6. Even though mass_function_msun is the strongest feature overall (with small values being associated with a classification as outlier) a small value alone is not sufficient for a classification as outlier as exemplified by Gaia DR3 246890014559489792 and in Figures 11 and 12.
These kinds of observations are interesting for complementing a human-based analysis of the data since one can use the models’ identification of important features to identify and investigate potential confounders and spurious relationships in the data. In some cases, such as for the eliminated features like parallax and absolute source magnitude, this is trivial but in others it can inspire follow-up investigations.
The most obvious example of this is the radial_velocity_error. The models consistently identified it as a strong feature and, as displayed in Figure 13, there is a correlation between the presence of a confirmed exoplanet or brown-dwarf companion and a smaller radial_velocity_error. As explained in Section 4.4.1, this particular feature has a complex (expected) relation with the presence of orbital signatures. All hosts of confirmed substellar companions have RV errors smaller than the median value for astrometric orbits (Fig. 13). The same applies to almost all our identified candidates. The underlying reason is expected to be a mixture of outlier classification (for small RV errors) and inlier classification (for large RV errors that can be caused by unmodelled binary motion). A detailed investigation of these effects is outside the scope of this initial proof of concept. We note, however, that we performed ablation studies that indicated that our approach identifies similar candidates without the radial_velocity_error feature. Thus, even if this relationship is spurious, it should not compromise the performance of the presented approach.
Figure 13: Distribution of radial velocity error in the sub-populations. The central peaks are caused by data imputation (Section 2.1) with the median for missing RV error values.
6 Conclusions
We successfully demonstrated a new approach for identifying individual systems with particular properties, i.e. outliers, in the ensemble of Gaia DR3 astrometric orbits. As first application, we searched for the most likely sources orbited by extrasolar planets and brown dwarfs, but the approach could be adapted to other science cases, e.g. to search for black holes or twin binary stars instead.
Using machine learning, we identified 22 high-confidence systems with substellar companions, of which five host extrasolar-planet candidates that were previously known. One of the exoplanet candidates (BD+24 4592, Gaia-ASOI-005) is a false-positive because the Gaia orbit does not correspond to the true orbit of the system, which is a known RV binary.
We identified 17 high-confidence systems with brown-dwarf-companion candidates. Two of those candidates (HD 104289 and HD 153376) are false positives because the Gaia orbits do not correspond to the true orbits of the known RV binaries, and one candidate (BD-06 2423A) has in fact a low-mass stellar companion although it was picked up by our models. Three brown-dwarf companion candidates (HD 206484, HD 78631, GSC 09436-0108) have combined astrometric+RV orbital solutions (‘AstroSpectroSB1’) and the companions straddle the substellar-stellar mass boundary. The companion of HD 106888 is a very low-mass star.
The presence of a brown dwarf with an astrometric mass of ${M_{2}}={62.43}^{+6.16}_{-5.44}\,{M_{\mathrm{Jup}}}$ orbiting our high-confidence candidate G 15-6 was confirmed by literature RV measurements. We propose that the remaining nine high-confidence brown-dwarf companion candidates identified in this study be followed-up preferentially.
With Gaia we anticipate that machine learning analyses of astrometric orbits will grow into becoming powerful complements to the traditional approaches of cross-checking with existing data and literature and following-up with additional observations when pursuing specific scientific goals.
As the Gaia astrometric-orbit catalogue will grow significantly in DR4, machine learning approaches will become more and more important for prioritising follow-up and identifying interesting systems and ‘outliers‘. In the same way, any analysis of the DR4 epoch data (Gaia astrometric timeseries of all sources will be released with DR4888https://www.cosmos.esa.int/web/gaia/release) that has the goal of identifying non-single stars, e.g. through non-linear motions and drifts, single-, or multi-companion orbits, will have to address the problem of false-positives. The holistic machine learning approach that facilitates the use of as much information as available has many advantages in that respect. In particular, once sufficient training data will be available, it is very likely that machine learning algorithms will automatically identify a sizeable fraction of twin binary stars, which are the principal astrophysical false positives that affect Gaia astrometric planet searches.
[…]
Acknowledgements
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France (Wenger et al., 2000). The authors made use of astropy (a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013), scipy (Jones et al., 2001), numpy (Oliphant, 2007), ipython (Pérez & Granger, 2007), pandas (Reback et al., 2022), matplotlib (Hunter, 2007), and astroquery (Ginsburg et al., 2019) This research made use of pystrometry, an open source Python package for astrometry timeseries analysis (Sahlmann, 2019). This project made use of the ESA Datalabs platform.
The authors acknowledge the positive role of the ESAC canteen and its friendly staff for facilitating informal exchanges during breakfast. This project is the outcome of one of these.
Data Availability
The Gaia data underlying this research are available at the Gaia archive at ESA (DOI: https://doi.org/10.5270/esa-qa4lep3). The relevant tables are gaiadr3.gaia_source, gaiadr3.nss_two_body_orbit, and gaiadr3.binary_masses.
Appendix A Auxiliary tables
Table 5: Used and computed fields for the gaia_source table. Instead of reproducing the field descriptions from van Leeuwen et al. (2022) here, we included a direct link to the online documentation.
Column Description Comment
source_id online documentation not used
ra online documentation
dec online documentation
parallax online documentation confounder
parallax_error online documentation
pmra online documentation confounder
pmdec online documentation confounder
visibility_periods_used online documentation
radial_velocity online documentation
radial_velocity_error online documentation
astrometric_chi2_al online documentation
ipd_gof_harmonic_amplitude online documentation
bp_rp online documentation
phot_bp_rp_excess_factor online documentation
ruwe online documentation
phot_g_mean_mag online documentation confounder
phot_rp_mean_mag online documentation
phot_bp_mean_mag online documentation
astrometric_excess_noise_sig online documentation
astrometric_primary_flag online documentation not used
astrometric_excess_noise online documentation
astrometric_n_good_obs_al online documentation
absolute_phot_g_mean_mag absolute $G$ mag computed
Table 6: Used and computed fields for the nss_two_body_orbit table.
Column Description Comment
source_id online documentation not used
nss_solution_type online documentation not used
ra online documentation
dec online documentation
parallax online documentation confounder
parallax_error online documentation
pmra online documentation confounder
pmdec online documentation confounder
period online documentation
period_error online documentation
t_periastron online documentation
t_periastron_error online documentation
eccentricity online documentation
eccentricity_error online documentation
astrometric_n_good_obs_al online documentation
obj_func online documentation
goodness_of_fit online documentation
efficiency online documentation
significance online documentation
p1_a0_mas $a_{0}$ (mas) computed
p1_omega_deg $\omega$ (deg) computed
p1_OMEGA_deg $\Omega$ (deg) computed
p1_incl_deg $i$ (deg) computed
mass_function_msun $f_{M}$ ($M_{\sun}$) computed
Table 7: Assigned labels
Gaia DR3 source_id Name label Reference
2047188847334279424 HD185501 binary_star Horch et al. (2020)
5122670101678217728 HD12357 binary_star Marcussen & Albrecht (2023)
2052469973468984192 Ross 1063 binary_star Marcussen & Albrecht (2023)
685029558383335168 HD77065 brown_dwarf_companion Gaia Collaboration et al. (2023b)
3751763647996317056 HD89707 brown_dwarf_companion Gaia Collaboration et al. (2023b)
3750881083756656128 HD91669 brown_dwarf_companion Gaia Collaboration et al. (2023b)
3309006602007842048 HD30246 brown_dwarf_companion Gaia Collaboration et al. (2023b)
2778298280881817984 HD5433 brown_dwarf_companion Gaia Collaboration et al. (2023b)
2651390587219807744 BD-004475 brown_dwarf_companion Gaia Collaboration et al. (2023b)
5563001178343925376 HD52756 brown_dwarf_companion Gaia Collaboration et al. (2023b)
824461960796102528 HD82460 brown_dwarf_companion Gaia Collaboration et al. (2023b)
43574131143039104 LHS1610 brown_dwarf_companion Fitzmaurice et al. (2023)
1035000055055287680 HD68638A brown_dwarf_companion Unger et al. (2023)
5999024986946599808 CD-4610046 brown_dwarf_companion Unger et al. (2023)
1181993180456516864 HD132032 brown_dwarf_companion Unger et al. (2023)
873616860770228352 BD+291539 brown_dwarf_companion Gaia Collaboration et al. (2023b)
855523714036230016 HD92320 brown_dwarf_companion Gaia Collaboration et al. (2023b)
6421118739093252224 HD175167 exoplanet Gaia Collaboration et al. (2023b)
3424193536079703808 HD39392 exoplanet Stevenson et al. (2023)
2603090003484152064 GJ876 exoplanet Gaia Collaboration et al. (2023b)
4976894960284258048 HD142 exoplanet Gaia Collaboration et al. (2023b)
637329067477530368 HD81040 exoplanet Gaia Collaboration et al. (2023b)
5855730584310531200 HD111232 exoplanet Gaia Collaboration et al. (2023b)
2367734656180397952 BD-170063 exoplanet Gaia Collaboration et al. (2023b)
4745373133284418816 HR810 exoplanet Gaia Collaboration et al. (2023b)
1594127865540229888 HD132406 exoplanet Gaia Collaboration et al. (2023b)
4062446910648807168 HD164604 exoplanet Gaia Collaboration et al. (2023b)
4698424845771339520 WD 0141-675 false_positive_orbit Cosmos Pages
5765846127180770432 HIP 64690 false_positive_orbit Cosmos Pages
5266148569447305600 HD42936 very_low_mass_stellar_companion Barnes et al. (2020)
1224551770875466496 HD140913 very_low_mass_stellar_companion Unger et al. (2023)
4753355209745022208 HD17155 very_low_mass_stellar_companion Unger et al. (2023)
Table 8: All 66 lower-confidence candidates selected as having $0.50.125<\rho<=$ 0.5. Colums are the same as in Table 4.
Gaia DR3 source_id Name SpT sol. $P$ $f_{M}$ $M_{2}$ $\rho$ $\rho_{\mathrm{ssc}}$ $\rho_{\mathrm{nss}}$ ssc $M_{1}$ $M_{2,\mathrm{alt}}$
type (day) ($M_{\sun}$) ($M_{\mathrm{Jup}}$) ($M_{\sun}$) ($M_{\sun}$)
932447162423519232 UCAC4 684-047911 M2.3 ASB1 407.6 1.2e-05 18.0 0.2 0.0 0.5 0 ${0.63}^{+0.05}_{-0.15}$ $0.150$-$0.639$
4632430611684633344 HD 14717 G0V Orbital 778.4 1.4e-05 27.2 0.2 0.5 0.0 1 ${1.10}^{+0.05}_{-0.10}$ $0.022$-$1.208$
4652393894560537728 HD 31251 G2V ASB1 419.8 8.1e-05 46.6 0.2 0.0 0.5 0 NaN NaN
2098419251579450880 2MASS J18405933+3950016 Orbital 241.7 8.9e-05 36.3 0.2 0.2 0.2 1 ${0.65}^{+0.16}_{-0.16}$ $0.029$-$0.936$
2984269003840680832 BD-15 1060 G5 Orbital 424.9 9.5e-05 52.2 0.2 0.5 0.0 1 ${1.09}^{+0.06}_{-0.09}$ $0.045$-$1.291$
3205578257602278784 TYC 4730-512-1 Orbital 598.1 9.7e-05 49.1 0.2 0.2 0.2 1 ${0.98}^{+0.06}_{-0.09}$ $0.041$-$1.171$
2994437527894182272 HD 42606 K3V OTS 800.1 9.8e-05 46.4 0.2 0.0 0.5 0 ${0.90}^{+0.05}_{-0.16}$ $0.516$-$0.984$
5981932494567274368 CD-51 9508 G8 ASB1 413.9 1.0e-04 48.7 0.2 0.5 0.0 1 ${0.96}^{+0.06}_{-0.06}$ ${0.044}^{+0.008}_{-0.008}$
4854976609869179648 CD-41 1115 G5 ASB1 406.5 1.2e-04 50.6 0.2 0.5 0.0 1 ${0.94}^{+0.06}_{-0.06}$ ${0.068}^{+0.008}_{-0.009}$
2217319511295868672 BD+62 1950 F8 Orbital 596.0 1.2e-04 60.7 0.4 0.5 0.2 1 ${1.20}^{+0.07}_{-0.14}$ $0.054$-$1.362$
4598289263814843136 4598289263814843136 N/A Orbital 435.8 1.4e-04 45.2 0.2 0.2 0.2 1 ${0.71}^{+0.05}_{-0.21}$ $0.040$-$0.726$
5671384265738137984 2MASSI J0952219-192431 M7Ve Orbital 278.5 1.5e-04 16.8 0.2 0.2 0.2 1 ${0.15}^{+0.05}_{-0.08}$ $0.046$-$0.226$
2744694491118490752 BD+05 5218 G0 ASB1 248.0 1.7e-04 59.4 0.2 0.5 0.0 1 ${0.98}^{+0.06}_{-0.06}$ ${0.042}^{+0.005}_{-0.005}$
1543883104727682176 TYC 3456-251-1 Orbital 593.3 1.7e-04 55.8 0.2 0.2 0.2 1 ${0.88}^{+0.05}_{-0.30}$ $0.050$-$0.827$
6604510475374443392 6604510475374443392 N/A Orbital 761.5 1.9e-04 57.2 0.2 0.2 0.2 1 ${0.88}^{+0.06}_{-0.14}$ $0.048$-$1.025$
5077267349557463680 TYC 6437-159-1 Orbital 436.1 2.0e-04 61.3 0.2 0.2 0.2 1 ${0.95}^{+0.06}_{-0.09}$ $0.054$-$1.168$
3087990814774129920 TYC 198-493-1 Orbital 722.6 2.0e-04 56.1 0.4 0.5 0.2 1 ${0.83}^{+0.05}_{-0.27}$ $0.047$-$0.812$
5488452086658908544 HD 65577 F0IV Orbital 678.3 2.1e-04 85.7 0.2 0.2 0.2 1 ${1.54}^{+0.05}_{-0.19}$ $0.072$-$1.719$
1856067303769195648 HD 340935 G2 Orbital 1176.2 2.1e-04 63.3 0.2 0.5 0.0 1 ${0.97}^{+0.06}_{-0.06}$ ${0.058}^{+0.006}_{-0.006}$
6500589824638636416 HD 219145 F5V Orbital 316.0 2.1e-04 73.9 0.2 0.5 0.0 1 ${1.21}^{+0.06}_{-0.19}$ $0.065$-$1.342$
766977465670727936 HD 100360 K0 OTS 572.1 2.3e-04 66.7 0.2 0.0 0.5 0 NaN NaN
4189636354105111168 TYC 5733-1115-1 ASB1 947.1 2.5e-04 68.5 0.2 0.0 0.5 0 NaN NaN
273199953420523648 TYC 3734-75-1 Orbital 391.5 2.6e-04 64.4 0.2 0.2 0.2 1 ${0.88}^{+0.05}_{-0.30}$ $0.056$-$0.854$
5144739739589028992 LP 769-41 Orbital 150.8 2.9e-04 47.2 0.2 0.2 0.2 1 ${0.52}^{+0.05}_{-0.17}$ $0.064$-$0.589$
4224582166524567040 4224582166524567040 N/A Orbital 413.7 2.9e-04 50.8 0.4 0.2 0.5 1 ${0.58}^{+0.05}_{-0.17}$ $0.045$-$0.657$
4668056781289612928 CD-69 203 ASB1 429.0 3.1e-04 79.8 0.2 0.5 0.0 1 ${1.11}^{+0.05}_{-0.11}$ $0.070$-$1.357$
6574387911224287488 L 499-75 M3 OTS 240.3 3.3e-04 50.0 0.5 0.5 0.5 1 ${0.53}^{+0.05}_{-0.18}$ $0.045$-$0.597$
1973932545092255616 TYC 3196-1933-1 Orbital 402.0 3.3e-04 81.5 0.4 0.5 0.2 1 ${1.11}^{+0.06}_{-0.09}$ $0.070$-$1.393$
5905528909017796480 5905528909017796480 N/A Orbital 606.6 3.4e-04 72.3 0.4 0.5 0.2 1 ${0.92}^{+0.06}_{-0.19}$ $0.062$-$1.050$
2717981134566374272 V* V335 Peg F5 OTS 787.7 3.5e-04 90.0 0.5 0.5 0.5 1 ${1.25}^{+0.06}_{-0.14}$ $0.079$-$1.500$
2347301645624022784 CD-25 215 Orbital 474.2 3.6e-04 78.4 0.2 0.2 0.2 1 ${1.00}^{+0.06}_{-0.17}$ $0.066$-$1.182$
2042732453670353664 2042732453670353664 N/A Orbital 497.0 3.6e-04 72.9 0.2 0.2 0.2 1 ${0.89}^{+0.06}_{-0.14}$ $0.062$-$1.077$
6717694064902244224 TYC 7922-716-1 ASB1 288.7 3.6e-04 68.5 0.5 0.5 0.5 1 ${0.81}^{+0.05}_{-0.06}$ ${0.067}^{+0.012}_{-0.012}$
4503350935885071872 TYC 1557-2079-1 Orbital 462.6 3.7e-04 82.6 0.4 0.5 0.2 1 ${1.07}^{+0.06}_{-0.10}$ $0.074$-$1.362$
5501729312336121856 UCAC2 9182345 ASB1 130.3 3.8e-04 62.2 0.5 0.5 0.5 1 ${0.68}^{+0.05}_{-0.05}$ ${0.061}^{+0.010}_{-0.010}$
2214819118774573440 BD+68 1377 G0 ASB1 310.9 4.2e-04 81.3 0.2 0.5 0.0 1 ${0.97}^{+0.06}_{-0.06}$ ${0.081}^{+0.009}_{-0.010}$
4474368187335444352 TYC 438-609-1 Orbital 508.1 4.3e-04 79.4 0.4 0.5 0.2 1 ${0.93}^{+0.05}_{-0.14}$ $0.068$-$1.143$
1378974777381950080 TYC 3061-368-1 Orbital 309.3 4.5e-04 77.2 0.2 0.5 0.0 1 ${0.87}^{+0.05}_{-0.30}$ $0.069$-$0.837$
6827680034091744896 HD 203771 K1V OTS 1105.4 4.5e-04 73.0 0.2 0.2 0.2 1 ${0.80}^{+0.05}_{-0.19}$ $0.065$-$0.920$
4984264505688458112 UCAC4 245-001421 ASB1 196.7 4.6e-04 80.0 0.5 0.5 0.5 1 ${0.91}^{+0.06}_{-0.06}$ ${0.075}^{+0.009}_{-0.010}$
6228661976111571072 HD 131197 K2III/IV Orbital 421.0 4.8e-04 79.0 0.4 0.2 0.5 1 ${0.87}^{+0.05}_{-0.15}$ $0.070$-$1.063$
5538678297515561088 CD-38 3568 ASB1 267.6 5.0e-04 87.4 0.5 0.5 0.5 1 ${1.00}^{+0.06}_{-0.07}$ ${0.088}^{+0.013}_{-0.014}$
4390684426062198528 TYC 413-602-1 ASB1 240.9 5.1e-04 86.5 0.2 0.5 0.0 1 ${0.96}^{+0.06}_{-0.07}$ $0.073$-$1.280$
1693282786205650560 2MASS J12114960+7752445 Orbital 651.2 5.2e-04 54.3 0.2 0.2 0.2 1 ${0.47}^{+0.05}_{-0.17}$ $0.048$-$0.548$
6186817449775754496 PM J13128-2603 ASB1 87.0 5.2e-04 64.2 0.2 0.2 0.2 1 ${0.60}^{+0.05}_{-0.05}$ ${0.075}^{+0.013}_{-0.013}$
6345830951689788672 CPD-85 499 G5 ASB1 485.0 5.7e-04 92.8 0.2 0.5 0.0 1 ${1.01}^{+0.06}_{-0.07}$ ${0.080}^{+0.012}_{-0.012}$
5829002521785164160 TYC 9040-2811-1 ASB1 508.3 6.4e-04 107.3 0.2 0.5 0.0 1 ${1.19}^{+0.06}_{-0.06}$ ${0.087}^{+0.014}_{-0.015}$
4188152322648398592 2MASS J19262435-1045398 Orbital 175.8 7.0e-04 49.8 0.2 0.2 0.2 1 ${0.34}^{+0.05}_{-0.16}$ $0.043$-$0.422$
548478735430284416 BD+74 118 K5V OTS 48.1 7.6e-04 86.9 0.4 0.5 0.2 1 ${0.78}^{+0.05}_{-0.21}$ $0.078$-$0.920$
5344309621263564160 TYC 8631-835-1 ASB1 789.9 7.7e-04 97.0 0.2 0.5 0.0 1 ${0.92}^{+0.05}_{-0.06}$ ${0.078}^{+0.009}_{-0.009}$
6012944872855490944 HD 139096 G3V ASB1 129.5 8.3e-04 104.6 0.2 0.2 0.2 1 ${1.00}^{+0.06}_{-0.06}$ ${0.079}^{+0.009}_{-0.009}$
4441168914767151360 HD 153402 K0 OTS 104.2 8.4e-04 96.1 0.2 0.0 0.5 0 ${0.87}^{+0.05}_{-0.19}$ $0.088$-$1.071$
3063883403860016512 TYC 4858-1465-1 ASB1 328.7 9.4e-04 102.4 0.2 0.0 0.5 0 ${0.90}^{+0.05}_{-0.05}$ ${0.101}^{+0.015}_{-0.016}$
4725142772167859840 L 127-40 ASB1 118.8 1.1e-03 96.2 0.2 0.2 0.2 1 ${0.74}^{+0.06}_{-0.05}$ ${0.084}^{+0.013}_{-0.012}$
3067929881523376384 TYC 4854-2258-1 ASB1 262.7 1.1e-03 117.0 0.2 0.0 0.5 0 ${0.99}^{+0.06}_{-0.06}$ ${0.182}^{+0.040}_{-0.038}$
1003812373374277248 HD 49039 G5 OTS 19.9 1.5e-03 117.4 0.2 0.0 0.5 0 ${0.87}^{+0.06}_{-0.21}$ $0.100$-$1.122$
4955501899980809088 TYC 7548-325-1 ASB1 337.3 1.8e-03 125.2 0.2 0.2 0.2 1 ${0.86}^{+0.06}_{-0.06}$ ${0.082}^{+0.013}_{-0.013}$
5247111556405603840 TYC 8952-212-1 ASB1 402.4 2.2e-03 126.2 0.2 0.5 0.0 1 ${0.76}^{+0.06}_{-0.05}$ ${0.098}^{+0.020}_{-0.019}$
5612492292661869824 HD 60666 K1III OTS 29.1 2.3e-03 150.3 0.2 0.0 0.5 0 NaN NaN
2743198605549406464 HD 223238 G5V OTS 61.3 2.5e-03 157.9 0.2 0.0 0.5 0 ${1.01}^{+0.06}_{-0.10}$ $0.133$-$1.364$
6911090115750857344 SCR J2108-0600 ASB1 144.1 2.9e-03 144.0 0.2 0.5 0.0 1 ${0.81}^{+0.05}_{-0.05}$ ${0.094}^{+0.015}_{-0.015}$
798068905726303232 * 11 LMi G8Va OTS 27.1 3.0e-03 156.2 0.2 0.0 0.5 0 ${0.91}^{+0.05}_{-0.11}$ $0.137$-$1.182$
6795834500861297408 SIPS J2049-2800 sdM6.5? Orbital 641.1 3.1e-03 52.0 0.2 0.2 0.2 1 ${0.15}^{+0.05}_{-0.07}$ $0.065$-$0.348$
3931991608291720192 TYC 880-843-1 ASB1 257.6 3.4e-03 184.5 0.2 0.5 0.0 1 ${1.10}^{+0.06}_{-0.06}$ ${0.093}^{+0.016}_{-0.015}$
845583819682861184 LSPM J1131+5627 Orbital 392.5 1.3e-02 84.1 0.2 0.2 0.2 1 ${0.12}^{+0.05}_{-0.05}$ ${0.080}^{+0.019}_{-0.018}$
5207515806222756480 CPD-81 193 ASB1 71.8 1.5e-02 292.0 0.2 0.5 0.0 1 ${0.92}^{+0.05}_{-0.05}$ ${0.080}^{+0.017}_{-0.016}$
Figure 14: Decision plot showing feature contribution to model output for the random forest models for the brown dwarf in HD 89707. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 15: Decision plot showing feature contribution to model output for the XGBoost models for the brown dwarf in HD 89707. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.